using -1 as a factor (divide)
(x-4)^2
__________ =
(4-x)^2


thanx

Answer by fractalier(2101)

(x-4)^2 / (4-x)^2 =
(x-4)^2 / (4-x(4-x) =
(x-4)^2 / (-1)(x-4)(-1)(x-4) =
(x-4)^2 / (x-4)^2 =
1